OpenStudy (anonymous):

How do I solve 20cos squared x+cos x-1=0 [0,pi]

hartnn (hartnn):

can you solve quadratics ? this is quadratic in cos x ....

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

a=20,b=1,c=-1?

hartnn (hartnn):

yes! cos x = -b+//sqrt b^2-4ac /2a.... i think you know the formula...

OpenStudy (anonymous):

I got -1/4 and 1/5(is that the end of the problem)?

hartnn (hartnn):

correct! so, x = cos inverse of -1/4 x=cos inverse of 1/5

hartnn (hartnn):

do you have choices/options?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

it is open-ended

OpenStudy (anonymous):

so would the answers be cos inverse=-1/4,1/5?

hartnn (hartnn):

yeah

OpenStudy (anonymous):

that's it?

hartnn (hartnn):

like this: x = cos inverse of -1/4 x=cos inverse of 1/5

OpenStudy (anonymous):

ty so much

hartnn (hartnn):

welcome :)